OpenStudy (anonymous):

2x/x-3 + 6/x+3= -28/x^2-9

jhonyy9 (jhonyy9):

2x/(x-3) +6/(x+3) =-28/(x*2 -9) --- (2x(x+3)+6(x-3))/(x*2 - 9) = -28/(x*2 -9) (2x*2 +6x+6x-18) = -28 --- 2x*2 +12x - 18 = -28 /2 --- x*2 +6x -9 = -14 x*2 +6x +5 =0